countryboy3642:

Which set of side lengths form a right triangle? 14 m, 20 m, 25 m 10 cm, 6 cm, 8 cm 3 ft, 6 ft, 5 ft 7 cm, 8 cm, 10 cm

harliii:

14*14+20*20=25*25

countryboy3642:

so its a

harliii:

no im not done help u yet

countryboy3642:

ok

harliii:

try 10*10+6*6=8*8

harliii:

then try 3*3+6*6=5*5

harliii:

7*7+8*8=10*10

countryboy3642:

i dont have time for this

harliii:

what ever answer is equal to the expression

countryboy3642:

just tell me the answer

harliii:

i cant give the answer

Hero:

@countryboy3642 are you familiar with the Pythagorean Theorem \(a^2 + b^2 = c^2\)? If so, then all you have to do is just plug each trio of side lengths into that formula while making sure that \(c\) is the longest side length you place into the formula. Start with the first set of given side lengths and keep trying each given trio until both sides equal. Once both sides are equal, then you will have found the correct trio.

Hero:

@countryboy3642 are you willing to try this? If not, then we can't help you. This site is not an answer-giving site. This site exists to help you understand how to solve these problem ON YOUR OWN. We're only here to help GUIDE YOU, but you must do the work yourself.

countryboy3642:

ok